INTERNAL MIXING, PHENYL RING TORSION AND EXCITONIC INTERACTION IN DIPHENYL-METHANE

NATHAN R. PILLSBURY, JAIME A. STEARNS, \underline{CHRISTIAN W. MÜLLER}, ALOKE DAS, TAL-ITHA M. SELBY and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-2084.

The close proximity of two identical ultraviolet chromophores render diphenylmethane (DPM) an interesting case for the study of the dependence of excitonic coupling on the distortion along low-frequency large-amplitude vibrational coordinates, in particular the phenyl ring torsional coordinates present in DPM. We have studied the fluorescence excitation spectrum and several single vibronic level fluorescence (SVLF) spectra of the Ã^{1}B(S_{1}) ← X̃^{1}A(S_{0}) and the B̃^{1}A(S_{2}) ← X̃^{1}A(S_{0}) transition of DPM cooled in a supersonic jet. The band in the excitation spectrum blue-shifted by 123 cm^{−1} from the S_{1} ← S_{0} origin was assigned to the S_{2} ← S_{0} origin. Its SVLF spectrum shows two-region fluorescence reminiscent of that previously observed in the gas-phase and supersonic jet spectra of naphthalene^{1} and ovalene^{2}. The high-energy region of this S_{2} 0^{0} SVLF spectrum was tentatively assigned to transitions originating in vibrationally excited levels of the S_{1} state. This finding indicates the presence of efficient internal mixing of the S_{2} 0^{0} level with the sparse manifold of S_{1} vibronic background levels. The quantum number changes in the non-totally symmetric torsional mode (see figure) upon internal mixing suggest that additional to the internal conversion transitions accounted for by the Jortner-Berry Δv = ±1 rule higher order vibronic mechanisms^{3} have to be considered as well.

(1) S. M. Beck, D. E. Powers, J. B. Hopkins and R. E. Smalley, J. Chem. Phys., 1980, 73, 2019.

(2) A. Amirav, U. Even and J. Jortner, J. Chem. Phys., 1981, 74, 3745.

(3) B. Scharf, Chem. Phys., 1975, 7, 478.